OpenStudy (anonymous):

simplify the rational expression 5x+10/x+2 x 2x/4x-10

OpenStudy (anonymous):

\[\frac{5x+10}{x+2} \times \frac{ 2x}{4x-10}=\frac{5(x+2)}{x+2} \times \frac{ 2x}{2(2x-5)}\] i.e. \[5 \times \frac{ x}{(2x-5)} = \frac{5 x}{(2x-5)} \]

OpenStudy (anonymous):

So the answer is 5x/2x-5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you could you answer another for me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Simplify the rational expression 3n^2-n/n^2-1 / n^2/n+1

OpenStudy (anonymous):

This can be expresssed as \[\frac{3n^2-n}{n^2-1} \times \frac{ n+1} {n^2}\] i.e. \[\frac{n(3n-)}{(n-1)(n+1)} \times \frac{ n+1} {n^2}\] Now cancel the like terms from the numerator as well as denominator, doing so you will get your ans

OpenStudy (anonymous):

Is it 1/n^2?

OpenStudy (anonymous):

Hi,are u there?

OpenStudy (sumi29):

You should get \[3n-1divn ^{2}-n\]

OpenStudy (anonymous):

is the answer 3n-1/n^2-n?

OpenStudy (sumi29):

Yes! You are right! :-)